OpenStudy (narissa):

4(x-3)+5x=-48 A.-4 B.all real numbers C.0 D.4

OpenStudy (narissa):

@vera_ewing

OpenStudy (vera_ewing):

So first, distribute 4(x-3) That would end up as: 4x-12

OpenStudy (vera_ewing):

So then, you have to isolate the variable, which is x.

OpenStudy (narissa):

why did you have to swich the 12 and the 5?

OpenStudy (vera_ewing):

4x - 12 + 5x = -48 4x + 5x - 12 = -48 4x + 5x - 12 + 12 = -48 +12 4x + 5x = -36 9x = -36 x = -4

OpenStudy (vera_ewing):

Ok, so I switched the 12 and the 5 because we have to get all of the x's on the right side.

OpenStudy (narissa):

ooh okay lol sorry im a bit slow lmao

OpenStudy (vera_ewing):

It's ok! :-)